OpenStudy (anonymous):

Divide 3/7 ÷ 1/2 * Please leave your answer as a fraction

geerky42 (geerky42):

To divide fraction, you just flip it then multiply. \(\dfrac{3}{7}\div\dfrac{1}{2}~\Longrightarrow~\dfrac{3}{7}\times\dfrac{2}{1}\) Is that clear?

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (anonymous):

yea -_-

jhonyy9 (jhonyy9):

so when divide twi fractions so multiplie the first by second invers

OpenStudy (anonymous):

im not a freakin reatard